OpenStudy (anonymous):

If you have 10.0 g of a substance that decays with a half-life of 14 days, then how much will you have after 42 days? A. 0.10 g B. 0.31 g C. 1.25 g D. 2.50 g

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large m\left( t \right) = {m_0}{e^{ - t/\tau }}\] where m_0 is the initial mass of our sample, m(t) is the mass of our sample at time t, and \tau is the half life

OpenStudy (anonymous):

ok! what do we plug in?

OpenStudy (michele_laino):

here is the next step: \[\Large m\left( {42} \right) = 10 \times {e^{ - 42/14}} = \frac{{10}}{{{e^3}}} = ...grams\]

OpenStudy (anonymous):

ok! what is e?

OpenStudy (anonymous):

oh wait sorry haha i totally blanked out :P

OpenStudy (anonymous):

we get 0.497870684?

OpenStudy (michele_laino):

e is the Neperus constant, namely: \[e = 2.71828\]

OpenStudy (anonymous):

yes:)

OpenStudy (michele_laino):

yes! I got 0.49

OpenStudy (michele_laino):

my formula is the correct one!

OpenStudy (anonymous):

yay!! what would the solution be though? 0.49 is not a choice :/

OpenStudy (michele_laino):

please wait, I'm checking my computation

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

I think that \[\tau \] is the average life, and it is given by the subsequent formula: \[\tau = \frac{{14}}{{0.693}} = 20.2\;days\]

OpenStudy (anonymous):

ohh okay! i am confused, though... which choice would be the accurate solution? :/

OpenStudy (michele_laino):

so the right answer is given by the subsequent computation: \[\Large m\left( {42} \right) = 10 \times {e^{ - 42/20.2}} = \frac{{10}}{{{e^{2.079}}}} = ...grams\]

OpenStudy (anonymous):

oh! ok! so we get 1.25! so choice C is the solution?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay! thank you!!:) okay! onto the next:)

OpenStudy (michele_laino):

:) ok!